Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Joint Operation and Readiness Logbook, Contact: T.Haas, A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 578     Entry time: 14 Nov 2018, 10:23
Author: Adriano Violante 
Location: General 
Severity: Information 
Subject: Status report after dispatch meeting 09.30 

SPB

  • Lopp 5 update - will be agreed between SPB and AE
  • Smaract limit switches issues - AE will have a look. In the meanwhile also the company of the controlles has  been asked by SPB
  • KB motor steps have different size with respect to encoders. This issue is known and a workaround / solution has to be discussed with AE offline
  • Pulse picker will be used and commissioning finalized.
  • Sample chamber encoder drift issue. Hasx to be investighated further. AE can support.

FXE

SQS

SCS

MID

HED

XPD

​​VAC

AE

  • MID Loop 11 - Fw received. Being tested.
  • PLCs for optical lasers in SASE3 and 2 control being set up
  • SASE3 support for commissioning ongoing 

ITDM

CAS

AoB

ELOG V3.1.4-7c3fd00